# Optimistic Online Non-stochastic Control via FTRL

### v1

@@ Abstract
This paper brings the concept of “optimism” to the new and promising framework of online Non-stochastic Control (NSC). Namely, we study how can NSC benefit from a prediction oracle of unknown quality responsible for forecasting future costs. The posed problem is first reduced to an optimistic learning with delayed feedback problem, which is handled through the Optimistic Follow the Regularized Leader (OFTRL) algorithmic family. This reduction enables the design of OptFTRL-C, the first Disturbance Action Controller (DAC) with optimistic policy regret bounds. These new bounds are commensurate with the oracle’s accuracy, ranging from $\mathcal{O}(1)$ for perfect predictions to the order-optimal $\mathcal{O}(\sqrt{T})$ even when all predictions fail. By addressing the challenge of incorporating untrusted predictions into control systems, our work contributes to the advancement of the NSC framework and paves the way towards effective and robust learning-based controllers.
@@ I Introduction
We study the NSC framework originally introduced in [1]: Consider a time-slotted dynamical system where at each time slot, the controller observes the system state $\boldsymbol{x}_{t}\in\mathbb{R}^{d_{x}}$ and decides an action $\boldsymbol{u}_{t}\in\mathbb{R}^{d_{u}}$ which then induces a cost $c_{t}(\boldsymbol{x}_{t},\boldsymbol{u}_{t})$, and causes a transition to a new state $\boldsymbol{x}_{t+1}$. Note that the cost and the new state are revealed to the controller after it commits its action. Similar to [1], we study Linear Time Invariant (LTI) systems where the state transition is parameterized by matrices $A\in\mathbb{R}^{d_{x}\times d_{x}}$, $B\in\mathbb{R}^{d_{x}\times d_{u}}$ and a disturbance vector $\boldsymbol{w}_{t}\in\mathbb{R}^{d_{x}}$:
@@
$\boldsymbol{w}_{t}$ can be arbitrarily set by an adversary, and we only restrict it to be upper-bounded, i.e., $\|\boldsymbol{w}_{t}\|\leq w,\forall t$. Similarly, the adversary is allowed to select any $l$-Lipschitz convex cost function $c_{t}:\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{u}}\mapsto\mathbb{R}$ at each slot. Since neither the cost nor disturbances are confined to follow a fixed and/or known distribution, the control problem is “Non-stochastic”.
The controller (or learner) aims to find a (possibly time-varying) policy that maps states to actions, $\pi:\boldsymbol{x}\mapsto\boldsymbol{u}$, from a given policy class $\Pi$, such that the cost trajectory $\{c_{t}(\boldsymbol{x}_{t},\boldsymbol{u}_{t})\}_{t=1}^{T}$ is as small as possible. To quantify the learner’s performance, we employ the policy regret metric, as presented in [2]. Intuitively, the policy regret measures the accumulated differences between the cost incurred by the learner’s policy, and that of a stationary unknown cost-minimizing policy designed with access to all future cost functions and disturbances:
@@
where $(\boldsymbol{x}_{t}(\pi),\boldsymbol{u}_{t}(\pi))$ is the counterfactual state-action sequence that would have emerged under the benchmark policy, whereas $(\boldsymbol{x}_{t},\boldsymbol{u}_{t})$ is the actual state-action sequence that emerged from following possibly different policies by the learner. A sub-linear regret means that the cost endured by the learner will converge to that of the optimal unknown policy at the same sub-linear rate, i.e., $\nicefrac{{\mathcal{R}_{T}}}{{T}}\rightarrow 0$ as $T\rightarrow\infty$. Note that the benchmark depends on the actual witnessed costs and disturbances, encoding a stronger concept of “adaptability” unlike the pessimistic $\mathcal{H}_{\infty}$ controllers, which assume worst case costs [3]. Remarkably, the Gradient Perturbation Controller (GPC) achieves $\mathcal{O}(T^{\nicefrac{{1}}{{2}}})$ regret, which is order optimal, and was shown to deliver superior performance to the more conventional $\mathcal{H}_{2}$ controllers [1].
@@
Unfortunately, such optimistic algorithms are yet to find their way to online NSC. This might be surprising given that previous NSC results were obtained through a streamlined reduction to the standard Online Convex Optimization (OCO) framework [8]. Hence, one might anticipate that optimistic NSC algorithms can be derived using a similar approach. Interestingly, this is not the case. Combining optimistic learning and NSC poses unique challenges for both frameworks. First, existing optimistic learning algorithms do not consider cost functions with memory and hence cannot handle states. Particularly, these algorithms update their regularizers at each time slot based on the prediction accuracy of the preceding slot’s cost [9, 10]. However, in stateful systems, an action made at $t$ will have an effect that spans across all slots until $t\!+\!d$, and thus uses predictions for all these slots. Since the accuracy of such multi-step predictions is not available until $t\!+d+\!1$, we cannot update the regularizer in the same standard way. Second, the guarantee of NSC is established via a reduction to the OCO with Memory (OCO-M) framework [11], which in turn is reduced to the standard OCO [12] via the concept of slowly moving decision variables [1, Thm. 4.6] [11, Thm. 3.1]. This later reduction cannot be utilized in optimistic learning where accurate predictions lead to little or no regularization, driving consecutive decisions of the optimistic algorithm to vary arbitrarily (up to the decision set diameter) [10, Sec $2.2$], [13, Sec $7.4$].
@@
This paper tackles exactly these challenges and aims to answer the question: is it possible to design an online learning algorithm whose policy regret is commensurate with the accuracy of an exogenous prediction oracle, while always staying sub-linear? Our paper answers this positively and builds upon recent advances in online learning, introducing, to our knowledge, the first optimistic controller for NSC.
@@
We achieve such optimistic guarantees by departing from the standard analysis approach of reducing the learner’s non-stationary policy to a stationary one [1, 14, 15], and instead directly analyzing the non-stationary policy. Specifically, to address the first challenge, we demonstrate that the additive separability of the linearized costs allows expressing the costs as a sum of memoryless but delayed functions of each of the decision variables. Next, to tackle the second challenge, we analyze the performance of each decision variable separately via an alternative reduction to the framework of “optimism with delay” [16]. Nonetheless, we customize this later framework with a specific “hint” design that exploits the structure of NSC where the cost is indeed delayed but still gradually being revealed at each step, leading to tighter bounds. We make these intuition-focused points concrete in our upcoming analysis.
@@
The main contribution is designing the first optimistic controller with policy regret that scales from $\mathcal{O}(1)$ to $\mathcal{O}(\sqrt{T})$, depending on the predictions’ accuracy. The methodology is based on a new perspective on stateful systems as systems with delayed feedback, for which we build upon recent results on delay and optimism. The next section reviews the related works. Sec. III provides the necessary background for our new algorithm, OptFTRL-C, introduced in Section IV. We then present numerical examples in Sec. V and conclude.
@@ II Related Work
The NSC problem was initiated in the seminal work of [1], which introduced the first controller with sub-linear policy regret for dynamical systems, generalizing the classical control problem to adversarial convex costs functions and adversarial disturbances. These results were further refined for strongly convex functions [17, 18, 19]; and systems where the actions are subject to fixed or adversarially-changing constraints [20, 21]. Follow-up works also looked at the NSC problem under more general assumptions on the system matrices, including unknown $(A,B)$ [2], systems with bandit feedback [22], and time-varying systems [23]. Expectedly, the regret bounds deteriorate in these cases, e.g., becoming $\mathcal{O}(T^{\nicefrac{{2}}{{3}}})$ for unknown systems and $\mathcal{O}(T^{\nicefrac{{3}}{{4}}})$ for systems with bandit feedback. Efficiency is also investigated in [15] with projection-free methods. Going beyond the typical bounds in terms of $T$, [24] introduced an adaptive FTRL based controller whose bound is proportional to the witnessed costs and perturbations: $\mathcal{O}((\sum_{t=1}^{T}g_{t}^{2})^{\nicefrac{{1}}{{2}}})$ instead of $\mathcal{O}(T^{\nicefrac{{1}}{{2}}})$, where $g_{t}$ depends on the witnessed costs and perturbations. These works do not consider the existence of an untrusted prediction oracle, as the model adopted here.
@@
Our work is also related to the robust MPC literature [30, 31], where (possibly inaccurate) predictions are used. However, the main difference is that while our algorithm is robust to bad predictions, it is not designed based on them. Specifically, our benchmark policy changes when worst case costs and predictions are not witnessed (as dictated by the regret metric). A newer line of research studies the MPC-style algorithms under the regret metric [32, 33]. Perfect predictions are assumed in [32], whose bound was later generalized in [33] with parameterized predictions error. While these works do not confine their policy class (i.e., directly optimizing $\{\boldsymbol{u}_{t}\}_{t}$), their regret is linearly dependent on the prediction error, which renders them suffer linear regret in the worst case scenario.
Optimism is an extension to the OCO framework that allows incorporating predictions of future costs without having the regret bound linearly depend on their quality. The optimistic learning framework, in its current prevalent form, originated in [9] where an Online Mirror Descent (OMD) algorithm was presented. Thereafter, optimism was studied under another equivalent111see [34, Sec. 6] for a discussion on the link between OMD and FTRL. framework (FTRL). Optimistic FTRL (OFTRL) has since undergone multiple improvements [13, 35] and we refer the reader to [36, Sec. 7.12] for comprehensive overview. Only recently, a close connection between delay and optimism emerged [16], a development we leverage in our current analysis of the NSC framework.
While optimistic learning has been studied for stochastic predictions [37, 38], and adversarial predictions [10, 5], these findings have not been applied to dynamical systems. In dynamical systems, the study of predictions is limited to either perfect predictions [39, 40], or fixed quadratic (thus, strongly convex) cost functions [41, 42, 43]. Our paper contributes towards filling this gap. Predictions may also be viewed via the lens of “context” [44] in the different problem of stochastic MDPs with finite state and action space. Lastly, the previous works of [43, 42] consider that the full predictions are provided a priori, meaning that the learner cannot benefit from updated (and possibly more accurate) predictions. Like MPC, we drop this assumption and allow the learner to use the most recent available predictions.
@@ III Preliminaries
DAC policy class. The class $\Pi$ under consideration in this paper, and in the broader NSC literature, is the Disturbance Action Controllers (DAC) (see [45, Ch. 6] for a general reference). The motivation behind DAC lies in the combination of expressive power and efficient parametrization. Specifically, DAC can approximate the large class of linear controllers, which, for instance, is guaranteed to include the universally optimal controller in the case of stochastic disturbances and quadratic costs (LQR settings). Simultaneously, the use of DAC actions has been demonstrated to induce convex cost functions in its parametrization.
A policy $\pi\in\Pi$, with a memory length of $p$, is parameterized by $p$ matrices $M\doteq\big{[}M^{[1]}|M^{[2]}|\dots|M^{[j]}|\dots|M^{[p]}\big{]}$, $M^{[j]}\in\mathbb{R}^{d_{u}\times d_{x}}$, and a fixed stabilizing controller $K$. We define the set $\mathcal{M}\doteq\{M:\|M\|\leq\kappa_{M}\}$, with the standard bounded variable assumption. The action at a step $t$ according to a policy $\pi_{t}\in\Pi$ is then calculated as follows:
@@
Note that $K$ is a stabilizing controller calculated a priori and provided as an input. This is formalized through the strong stability assumption [46, Def. 3.1], which is standard in NSC. It ensures that the DAC controller is supplied with a controller $K$ such that $\|(A+BK)^{t}\|_{\text{op}}\leq\kappa(1-\delta)^{t}$ for $\delta\in(0,1]$, where $\kappa>0$. Here, we make a slightly stricter assumption, enabling us to simplify the analysis:
@@
where we defined $\boldsymbol{\overline{w}}_{t-i-1}\doteq(\boldsymbol{w}_{t-i-1},\dots,\boldsymbol{w}_{t-i-p})$ so as to express the vector $\sum_{j=1}^{p}M_{t}^{[j]}\boldsymbol{w}_{t-j}$ compactly as $M_{t}\overline{\boldsymbol{w}}_{t-1}$. The above expression for the state can be obtained by simply unrolling the dynamic [^2] Assuming that the initial state $\boldsymbol{x}_{1}$ (before executing $M_{1},\dots,M_{t}$) is $\boldsymbol{0}$. in (1). This is proven, e.g., in [45, Lem. 7.3] and [25, Lem. 4.3].
Cost functions. While our guarantees continue to hold for the general convex cost function, we assume here that the cost functions are linearized:
@@
The cost is linear in the state and action $c_{t}(\boldsymbol{x}_{t},\boldsymbol{u}_{t})=\langle{\boldsymbol{\alpha}_{t}},{\boldsymbol{x}_{t}}\rangle+\langle{\boldsymbol{\beta}_{t}},{\boldsymbol{u_{t}}}\rangle.$
The linearity assumption provides a useful structure in our analysis (separability) and enables us to quantify the prediction error in terms of the parameters $\boldsymbol{\alpha}_{t}$ and $\boldsymbol{\beta}_{t}$. It does not, however, compromise the presented regret guarantees. Specifically, our bounds continue to hold for the general convex case. In fact, the linear costs are the most challenging [^3] As indicated in the preceding section, in the case of strongly convex costs, a refined and much tighter regret bound of $\mathcal{O}(\log(T))$ (compared to $\mathcal{O}(\sqrt{T})$) is possible [19]. in the online learning settings, and hence the regret caused by a sequence of general convex function can be indeed upper bounded by the regret caused by linearization of those functions. Thus, online learning works often focused on the linear case (see discussion on linearization in [34, Sec. 2.1] and [12, Sec. 2.4]). Future costs $\{c_{\tau}(\cdot,\cdot)\}_{\tau=t}^{t+d}$ can thus be predicted through the oracle output $\{\boldsymbol{\alpha}_{\tau},\boldsymbol{\beta}_{\tau}\}_{\tau=t}^{t+d}$.
Predictions. The prediction model we use has several advantages over those that appear in the preceding section. $(i)$ We allow the prediction oracle to update its forecast at every decision slot $t$ (similar to MPC). This flexibility is important, as in practice predictions can improve with time. $(ii)$ our analysis reveals that the parameter $d$ only needs to scale logarithmically with $T$. This implies that predictions are not required for the entire future. In fact, the proposed algorithm accommodates any number of predictions by simply setting $\boldsymbol{\alpha}_{\tau},\boldsymbol{\beta}_{\tau}$ to zero for steps where no prediction is available. In other words, our algorithm can operate with no predictions (an oracle that produces $\boldsymbol{0}$ for all $t$) or limited horizon predictions. $(iii)$ the presented algorithm and its guarantees place no assumptions on the predictions’ quality. To the best of our knowledge, this represents the most general prediction-based setting for online control.

### v2

@@ Abstract
This paper brings the concept of “optimism” to the new and promising framework of online Non-stochastic Control (NSC). Namely, we study how NSC can benefit from a prediction oracle of unknown quality responsible for forecasting future costs. The posed problem is first reduced to an optimistic learning with delayed feedback problem, which is handled through the Optimistic Follow the Regularized Leader (OFTRL) algorithmic family. This reduction enables the design of OptFTRL-C, the first Disturbance Action Controller (DAC) with optimistic policy regret bounds. These new bounds are commensurate with the oracle’s accuracy, ranging from $\mathcal{O}(1)$ for perfect predictions to the order-optimal $\mathcal{O}(\sqrt{T})$ even when all predictions fail. By addressing the challenge of incorporating untrusted predictions into online control, this work contributes to the advancement of the NSC framework and paves the way toward effective and robust learning-based controllers.
@@ I Introduction
We study the NSC framework originally introduced in [1]: Consider a discrete-time dynamical system where at each time slot, the controller observes the system state $\boldsymbol{x}_{t}\in\mathbb{R}^{d_{x}}$ and decides an action $\boldsymbol{u}_{t}\in\mathbb{R}^{d_{u}}$ which then induces a stage cost $c_{t}(\boldsymbol{x}_{t},\boldsymbol{u}_{t})$, and causes a transition to a new state $\boldsymbol{x}_{t+1}$. Note that the cost and the new state are revealed to the controller after it commits its action. Similar to [1], we study Linear Time Invariant (LTI) systems where the state transition is parameterized by matrices $A\in\mathbb{R}^{d_{x}\times d_{x}}$, $B\in\mathbb{R}^{d_{x}\times d_{u}}$ and a disturbance vector $\boldsymbol{w}_{t}\in\mathbb{R}^{d_{x}}$:
@@
$\boldsymbol{w}_{t}$ can be arbitrarily set by an adversary, and we only restrict it to be bounded, i.e., $\|\boldsymbol{w}_{t}\|\leq w,\forall t$. Similarly, the adversary is allowed to select any $l$-Lipschitz convex cost function $c_{t}:\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{u}}\mapsto\mathbb{R}$ at each slot. Since neither the cost nor disturbances are confined to follow a fixed and/or known distribution, the control problem is “Non-stochastic”.
The controller (or learner) aims to find a (possibly time-varying) policy that maps states to actions, $\pi:\boldsymbol{x}\mapsto\boldsymbol{u}$, from a given policy class $\Pi$, such that the stage costs $\{c_{t}(\boldsymbol{x}_{t},\boldsymbol{u}_{t})\}_{t=1}^{T}$ are as small as possible. To quantify the learner’s performance, we employ the policy regret metric, as presented in [2]. Intuitively, the policy regret measures the accumulated differences between the cost incurred by the learner’s policy, and that of a stationary unknown cost-minimizing policy designed with access to all future cost functions and disturbances:
@@
where $(\boldsymbol{x}_{t}(\pi),\boldsymbol{u}_{t}(\pi))$ is the counterfactual state-action sequence that would have emerged under the benchmark policy, whereas $(\boldsymbol{x}_{t},\boldsymbol{u}_{t})$ is the actual state-action sequence that emerged from following possibly different policies by the learner. A sub-linear regret means that the cost endured by the learner will converge to that of the optimal unknown policy at the same sub-linear rate, i.e., $\nicefrac{{\mathcal{R}_{T}}}{{T}}\rightarrow 0$ as $T\rightarrow\infty$. Note that the benchmark depends on the actual witnessed costs and disturbances, encoding a stronger concept of “adaptability” unlike the pessimistic $\mathcal{H}_{\infty}$ controllers, which assume worst case costs [3]. Remarkably, the Gradient Perturbation Controller (GPC) achieves $\mathcal{O}(\sqrt{T})$ regret, which is order optimal, and was shown to deliver superior performance to the more conventional $\mathcal{H}_{2}$ controllers [1].
@@ I-A Motivation
Unfortunately, such optimistic algorithms are yet to find their way to online NSC. This might be surprising given that previous NSC results were obtained through a streamlined reduction to the standard Online Convex Optimization (OCO) framework [8]. Hence, one might anticipate that optimistic NSC algorithms can be derived using a similar approach. Interestingly, this is not the case. Combining optimistic learning and NSC poses unique challenges for both frameworks. First, existing optimistic learning algorithms do not consider cost functions with memory and hence cannot handle states. Particularly, these algorithms update their regularizers at each time slot based on the accuracy of the prediction used by the preceding action [9, 10]. However, in stateful systems, an action made at $t$ will have an effect that spans across all slots until $t\!+\!d$, and thus uses predictions for all these slots. Since the accuracy of such multi-step predictions is not available until $t\!+d+\!1$, we cannot update the regularizer in the standard way. Second, the guarantee of NSC is established via a reduction to the OCO with Memory (OCO-M) framework [11], which in turn is reduced to the standard OCO [12] via the concept of slowly moving decision variables [1, Thm. 4.6] [11, Thm. 3.1]. This later reduction cannot be utilized in optimistic learning where accurate predictions lead to little or no regularization, driving consecutive decisions of the optimistic algorithm to vary arbitrarily (up to the set diameter) [10, Sec $2.2$], [13, Sec $7.4$].
@@
This paper tackles exactly these challenges and aims to answer the question: is it possible to design an online learning algorithm whose policy regret is commensurate with the accuracy of an exogenous prediction oracle, while always staying sub-linear? We answer this positively and builds upon recent advances in online learning, introducing, to our knowledge, the first optimistic controller for NSC.
I-B Contributions
@@
We achieve such optimistic guarantees by departing from the standard analysis approach of reducing the learner’s non-stationary policy to a stationary one [1, 14, 15], and instead directly analyzing the non-stationary policy. Specifically, to address the first challenge, we demonstrate that the additive separability of the linearized costs allows expressing the costs as a sum of memoryless but delayed functions of each of the decision variables. Next, to tackle the second challenge, we analyze the performance of each decision variable separately via an alternative reduction to the framework of “optimism with delay” [16]. Nonetheless, we customize this later framework with a specific “hint” design that exploits the structure of NSC where the cost is indeed delayed but still gradually being revealed at each step, leading to tighter bounds. We make these intuition-focused points concrete in the upcoming analysis.
@@
The main contribution is an optimistic controller with policy regret scaling from $\mathcal{O}(1)$ to $\mathcal{O}(\sqrt{T})$, based on prediction accuracy. The methodology hinges on a new perspective on stateful systems (with a prediction oracle) as systems with delayed feedback, for which we build upon recent results on delay and optimism. The next section reviews the related works. Sec. III provides the necessary background for the new algorithm, OptFTRL-C, introduced in Section IV. We then present numerical examples in Sec. V and conclude.
@@ II Related Work
The NSC problem was initiated in the seminal work of [1], which introduced the first controller with sub-linear policy regret for dynamical systems, generalizing the classical control problem to adversarial convex costs functions and adversarial disturbances. These results were further refined for strongly convex functions [17, 18, 19]; and systems with fixed or adversarially-changing constraints on the actions [20, 21]. Follow-up works also looked at the NSC problem under more general assumptions on the system matrices, including unknown $(A,B)$ [2], systems with bandit feedback [22], and time-varying systems [23]. Expectedly, the regret bounds worsen in these scenarios, such as increasing to $\mathcal{O}(T^{\nicefrac{{2}}{{3}}})$ for unknown systems and $\mathcal{O}(T^{\nicefrac{{3}}{{4}}})$ for systems with bandit feedback. Efficiency is also investigated in [15] with projection-free methods. Going beyond the typical bounds in terms of $T$, [24] introduced an adaptive FTRL based controller whose bound is proportional to the witnessed costs and perturbations: $\mathcal{O}((\sum_{t=1}^{T}g_{t}^{2})^{\nicefrac{{1}}{{2}}})$ instead of $\mathcal{O}({\sqrt{T}})$, where $g_{t}$ depends on the witnessed costs and perturbations. These works do not consider the existence of an untrusted prediction oracle, as the model adopted here.
@@
Our work is also related to the robust MPC literature [30, 31], where (possibly inaccurate) predictions are used. However, the main difference is that while our algorithm is robust to bad predictions, it is not designed based on them. Specifically, our benchmark policy changes when worst case costs and predictions are not witnessed (as dictated by the regret metric). A newer line of research studies the MPC-style algorithms under the regret metric [32, 33]. Perfect predictions are assumed in [32], whose bound was later generalized in [33] with parameterized predictions error. While these works consider the refined dynamic regret metric, their regret scales linearly with the prediction error.
Optimism extends the OCO framework by incorporating predictions of future costs, akin to “certainty equivalent” control. However, a key focus of optimistic algorithms is ensuring that the regret bound does not depend linearly on the accuracy of these predictions. The optimistic learning framework, in its current prevalent form, originated in [9] where an Online Mirror Descent (OMD) algorithm was presented. Thereafter, optimism was studied under another equivalent framework (FTRL). Optimistic FTRL (OFTRL) has since undergone multiple improvements [13, 34] and we refer the reader to [35, Sec. 7.12] for comprehensive overview. Only recently, a close connection between delay and optimism emerged [16], a development we leverage in our current analysis of the NSC framework.
While optimism has been studied for stochastic predictions [36, 37], and adversarial predictions [10, 5], these findings have not been applied to dynamical systems. In dynamical systems, the study of predictions is limited to either perfect predictions [38, 39], or fixed quadratic (thus, strongly convex) cost functions [40, 41, 42]. Our paper contributes to filling this gap. Predictions may also be viewed via the lens of “context” [43] in the problem of stochastic MDPs with finite states and actions. Lastly, the previous works of [42, 41] consider that the full predictions are provided a priori, meaning that the learner cannot benefit from updated predictions. Like MPC, we drop this assumption and allow the learner to use the most recent predictions.
@@ III Preliminaries
DAC policy class. The class $\Pi$ under consideration in this paper, and in the broader NSC literature, is the Disturbance Action Controllers (DAC) (see [44, Ch. 6] for a general reference). The motivation behind DAC lies in the combination of expressive power and efficient parametrization. Specifically, DAC can approximate the large class of linear controllers, which, for instance, is guaranteed to include the universally optimal controller in the LQR settings, and the linear positive systems with linear objective function [45]. Simultaneously, the use of DAC actions has been demonstrated to induce convex cost functions in its parametrization.
A policy $\pi\in\Pi$, with a memory length of $p$, is parameterized by $p$ matrices $M\doteq\big{[}M^{[1]}|M^{[2]}|\dots|M^{[j]}|\dots|M^{[p]}\big{]}$, $M^{[j]}\in\mathbb{R}^{d_{u}\times d_{x}}$, and a fixed stabilizing controller $K$. We define the set $\mathcal{M}\doteq\{M:\|M\|\leq\kappa_{M}\}$, with the bounded variable assumption. The action at a step $t$ according to a policy $\pi_{t}\in\Pi$ is then calculated as follows:
@@
$K$ is pre-calculated and provided as an input. This is formalized through the strong stability assumption [46, Def. 3.1], which is standard in NSC. It ensures the existence a controller $K$ such that $\|(A+BK)^{t}\|_{\text{op}}\leq\kappa(1-\delta)^{t}$ for $\delta\in(0,1]$, where $\kappa>0$. Here, we make a slightly stricter assumption, enabling us to simplify the analysis:
@@
where we defined $\boldsymbol{\overline{w}}_{t-i-1}\doteq(\boldsymbol{w}_{t-i-1},\dots,\boldsymbol{w}_{t-i-p})$ so as to express the vector $\sum_{j=1}^{p}M_{t}^{[j]}\boldsymbol{w}_{t-j}$ compactly as $M_{t}\overline{\boldsymbol{w}}_{t-1}$. The above expression for the state can be obtained by simply unrolling the dynamic in (1) and assuming, w.l.o.g that the initial state $\boldsymbol{x}_{1}$ (before executing $M_{1},\dots,M_{t}$) is $\boldsymbol{0}$. This is proven, e.g., in [44, Lem. 7.3] and [25, Lem. 4.3].
Cost functions. While our guarantees still hold for general convex cost functions, we focus here on the linear case:
@@
The cost is linear in the state and action $c_{t}(\boldsymbol{x}_{t},\boldsymbol{u}_{t})=\langle{\boldsymbol{\alpha}_{t}},{\boldsymbol{x}_{t}}\rangle+\langle{\boldsymbol{\beta}_{t}},{\boldsymbol{u_{t}}}\rangle.$ $\|\boldsymbol{\alpha}_{t}\|\leq\alpha$ and $\|\boldsymbol{\beta}\|\leq\beta$.
The linearity assumption provides a useful structure in the analysis (separability) and enables us to quantify the prediction error in terms of the parameters $\boldsymbol{\alpha}_{t}$ and $\boldsymbol{\beta}_{t}$. It does not, however, compromise the presented regret guarantees. In fact, the linear costs are the most challenging [^1] As indicated in Sec. II, in the case of strongly convex costs, a tighter regret bound of $\mathcal{O}(\log(T))$ (compared to $\mathcal{O}(\sqrt{T})$) is possible [19]. in the online learning settings; the regret caused by a sequence of general convex function can be indeed upper bounded by the regret caused by linearization of those functions. Thus, online learning works often focused on the linear case (see discussion on linearization in [12, Sec. 2.4]). Future costs can thus be predicted through the oracle’s output $\{\boldsymbol{\alpha}_{\tau},\boldsymbol{\beta}_{\tau}\}_{\tau=t}^{t+d}$.
Predictions. The prediction model we use has several advantages over those that appear in the preceding section. $(i)$ We allow the prediction oracle to update its forecast at every decision slot $t$ (similar to MPC). This flexibility is important, as in practice predictions can improve with time. $(ii)$ the analysis reveals that the parameter $d$ needs to scale logarithmically with $T$. This implies that predictions are not required for the entire future. $(iii)$ the presented algorithm and its guarantees place no assumptions on the predictions’ quality. To our knowledge, this represents the most general prediction-based setting for online control.